a fair dice is thrown twice. what is the probability that a 2 was scored on the first throw and a 4 on the second throw?
1
Expert's answer
2013-03-13T08:29:06-0400
Probability, that a 2 was scored on the first throw equals:
P1 = 1/6 (becausea dice has 6 angles )
Probability, that a4 was scored on the second throw equals:
P2 = 1/6
Finally, totalprobability equals:
P = P1*P2 = 1/6 * 1/6 =1/36
Answer: 1/36